Project

General

Profile

Actions

2017-12-05-physics-shifter » History » Revision 3

« Previous | Revision 3/14 (diff) | Next »
Rogers, Chris, 08 December 2017 12:25


2017-12-05-physics-shifter

2017-11-30 physics-shifter

Start Date: 2017-12-05
End Date:
Start Run: 10432
End Run:
Physics shifter: Chris Rogers

Summary

The aim was to get the empty data for corresponding LiH full settings:

  • 5th December - possibly recover magnets following a quench
  • 6th December - 2016-05.1 setting (flip mode, 140 MeV/c)
  • 7th December - 2017-02-7 setting (flip mode, 140 MeV/c)
  • 8th December - 2017-02-8 setting (flip mode, 170 MeV/c)

Elog, CDB

Online reco

Analysis

5th December

No data

6th December

Updated by Rogers, Chris over 6 years ago · 3 revisions